Solve the following pairs of simultaneous equations.
3x + 4y = -3
x + y = 1

Answer by mananth(16946)

3 x + 4 y = -3 .............1
Total value
1 x + 1 y = 1 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -4
3 x 4 y = -3
-4 x -4 y = -4
Add the two equations
-1 x = -7
/ -1
x = 7
plug value of x in (1)
3 x + 4 y = -3
21 + 4 y = -3
4 y = -3 -21
4 y = -24
y = -6
x=7,y=-6